A jeweler maker has 20 ounces of a gold alloy worth $160 an ounce and 10 ounce of a silver alloy worth $140 an ounce. How many ounces of another alloy at $124 an ounce must she add to these in order to have a mixture of the three that she can use to make 3-ounce jewelry worth $438 a piece?
And I really need to know at $124 an ounce answer. Thank you

Answer by ankor@dixie-net.com(22740)

A jeweler maker has 20 ounces of a gold alloy worth $160 an ounce and 10 ounce of a silver alloy worth $140 an ounce.
How many ounces of another alloy at $124 an ounce must she add to these in order to have a mixture of the three that she can use to make 3-ounce jewelry worth $438 a piece?
:
The 3 oz item worth $438; 438/3 = $146 an oz is the value of the resulting alloy
:
let x = amt of $124 alloy required
160(20) + 140(10) + 124x = 146(20+10+x)
3200 + 1400 + 124x = 146(x+30)
4600 + 124x = 146x + 4380
4600 - 4380 = 146x - 124x
220 = 22x
x = 220/22
x = 10 oz of $124 alloy required